OpenStudy (anonymous):

Short answer. Add the following:(-1.5x^2+7.9x+8.6)+(-5.5x^2-7.8x+5.2) this is the same problem except instead of this being -15 it is suppose to be -1.5, can you correct this for me and help me solve this mellylol

OpenStudy (saifoo.khan):

-1.5x^2 - 5.5x^2 = -7x^2

OpenStudy (anonymous):

\[-7. x^2+0.1 x+13.8 \]or\[\frac{1}{10} \left(-70 x^2+x+138\right) \]